# Tell Me What You See: An Iterative Deep Learning Framework for Image Captioning

### v1

@@ Abstract
Image captioning, a task at the confluence of computer vision and natural language processing, requires a sophisticated understanding of both visual scenes and linguistic structure. While modern approaches are dominated by large-scale Transformer architectures, this paper documents a systematic, iterative development of foundational image captioning models, progressing from a simple CNN-LSTM encoder-decoder to a competitive attention-based system. We present a series of five models, beginning with Genesis and concluding with Nexus, an advanced model featuring an EfficientNetV2B3 backbone and a dynamic attention mechanism. Our experiments chart the impact of architectural enhancements and demonstrate a key finding within the classic CNN-LSTM paradigm: merely upgrading the visual backbone without a corresponding attention mechanism can degrade performance, as the single-vector bottleneck cannot transmit the richer visual detail. This insight validates the architectural shift to attention. Trained on the MS COCO 2017 dataset, our final model, Nexus, achieves a BLEU-4 score of 31.4, surpassing several foundational benchmarks and validating our iterative design process. This work provides a clear, replicable blueprint for understanding the core architectural principles that underpin modern vision-language tasks.
Keywords: Image Captioning, Multimodal Learning, Deep Learning, Attention Mechanism, Information Bottleneck, Encoder-Decoder, CNN, RNN, LSTM, Spatial Encoder, MS COCO
@@ 1 Introduction
In this paper, we present the results of a structured, iterative research process aimed at building and refining an image captioning model. Our methodology is rooted in progressive enhancement, starting with a simple baseline and methodically introducing more complex components to address specific limitations. We document the development of five distinct models, named to reflect their role in this evolution: Genesis, Contexta, Clarity, Focalis, and Nexus.
This journey provides valuable insights into the interplay between visual perception, language modeling, and attention. We demonstrate empirically how each architectural modification impacts performance, including a counter-intuitive but instructive finding where a more powerful visual encoder led to a decrease in caption quality, thereby motivating the shift to an attention-based design. This work also serves as a pedagogical guide for students and practitioners, offering a clear, replicable framework to understand the architectural trade-offs that shaped modern vision-language models. Our final model, Nexus, leverages the lessons learned to achieve results that are competitive with, and on several key metrics surpass, foundational benchmarks in the field.
@@ 2 Related Work
The field of image captioning has evolved significantly since its inception. This section reviews the foundational encoder-decoder paradigms that our work builds upon, as well as the modern Transformer-based architectures that now define the state of the art.
@@ 2.1 Foundational Encoder-Decoder Architectures
Early work established the now-classic encoder-decoder architecture. Vinyals et al. [5] in their ”Show and Tell” model, successfully repurposed techniques from machine translation, using a pre-trained CNN like InceptionV3 as an encoder and an LSTM network as a decoder. Our initial model, Genesis, is a direct implementation of this foundational paradigm.
The primary limitation of this approach was the information bottleneck created by compressing the entire image into a single static context vector. The breakthrough came with the introduction of visual attention mechanisms by Xu et al. [6] in ”Show, Attend and Tell”. Their model allowed the decoder to dynamically ”look” at different parts of the image at each step of the caption generation process. This concept of soft-attention is central to our advanced models, Focalis and Nexus, and directly addresses the static bottleneck problem we investigate. Further refinements on this grid-based attention include the ”Bottom-Up and Top-Down Attention” model by Anderson et al. [7], which first identifies salient object regions and then applies top-down attention.
@@ 2.2 The Transformer Revolution and Modern Approaches
While attention-based Transformers are dominant, research into non-attention or attention-free architectures continues. For example, RWKV [24], a type of Recurrent Neural Network, has been adapted for multimodal tasks, offering a linear-time complexity alternative to the quadratic complexity of standard self-attention. This demonstrates that architectural diversity persists. Unlike the CNN-LSTM models with global average pooling that we investigate, architectures like RWKV or vision-specific designs like U-Nets [4] use different feature processing methods (recurrent states or skip connections) that may mitigate the specific single-vector bottleneck we identify.
Our work provides a crucial, hands-on analysis of the architectural trade-offs, particularly the attention-based resolution of the static information bottleneck, that motivated the field’s transition towards these more complex and dynamic systems.
@@ 2.3 Model 1: Genesis
The Genesis model serves as our foundational baseline, embodying the classic encoder-decoder paradigm that set the stage for modern image captioning. The architecture is intentionally simple to establish a clear performance benchmark. The encoder’s role is to ”see” the image, for which we employed the InceptionV3 architecture pre-trained on ImageNet. This choice provides a strong set of visual features learned from a massive dataset. The final convolutional feature map from InceptionV3 is passed through a Global Average Pooling layer, which condenses the entire spatial grid into a single, fixed-length 2048-dimension feature vector.
@@ 2.4 Model 2: Contexta
While Genesis established our visual baseline, its primary limitation was its decoder’s simplistic understanding of linguistic structure. A standard unidirectional LSTM only considers past context (the words it has already generated) when predicting the next word. To address this, we developed Contexta, a model designed to test the hypothesis that a richer language model could improve caption quality, even with the same static visual input.
@@ 2.5 Model 3: Clarity
Our next step was to test a straightforward hypothesis: would a more powerful visual feature extractor directly lead to better captions? To investigate this, we designed the Clarity model. We replaced the InceptionV3 backbone with a state-of-the-art EfficientNetV2B3, known for its superior accuracy. Critically, we retained the rest of the Contexta architecture, including the global average pooling encoder and the Bi-LSTM decoder.
@@ 2.6 Model 4: Focalis
The performance degradation of Clarity provided a critical diagnostic insight: a more powerful visual feature extractor is useless if its output cannot be effectively utilized by the decoder. This confirmed our hypothesis that the single, static feature vector was acting as a severe information bottleneck. Although EfficientNetV2B3 was providing a richer, more detailed map of the image, global average pooling collapsed this entire map into a single vector, irretrievably discarding the spatial information that defines ”where” objects are. To solve this, we designed Focalis, which represents a fundamental architectural pivot from a static to a dynamic vision-language interface.
The core innovation in Focalis is the introduction of a Luong-style attention mechanism, which completely removes the global average pooling layer. Instead of forcing the model to ”remember” the entire image from a single vector, we allow it to selectively ”look” at different regions of the image at each step of the caption generation process. This enables the decoder to ground specific words in specific spatial locations, for example, focusing on the region containing a ”dog” when generating that word, and then shifting focus to the ”ball” when describing what the dog is chasing.
@@
  • Feature Grid: The EfficientNetV2B3 encoder outputs a (10×10, 1536) feature grid, preserving the spatial layout of the image. After our spatial encoder (described below), this becomes a set of 100 context-aware feature vectors, which serve as the ”values” for the attention mechanism.
@@
Crucially, before the attention mechanism is applied, we introduce our key architectural refinement: a Bidirectional LSTM as a spatial encoder. Unlike standard attention mechanisms that treat each image patch independently, this Bi-LSTM processes the 100 patch vectors as a sequence. By reading the sequence of patches forwards and backwards, it enriches each patch’s vector with information about its neighbors. For instance, the vector for a patch containing a ”face” becomes aware that it is spatially adjacent to a patch containing ”hair.” This creates a more contextually rich set of feature vectors over which the decoder can attend.
@@
This complete architecture directly addresses the limitations observed in our previous models. By providing the decoder with dynamic, selective access to context-aware spatial features, we hypothesized that Focalis would not only resolve the performance degradation seen in Clarity but would significantly surpass all prior models. As the results show, this proved to be the case, marking the critical turning point in our development cycle.
@@ 2.7 Model 5: Nexus
The Nexus model does not introduce a new architecture; rather, it represents the culmination of our research by testing the scalability and full potential of our best design. While Focalis proved that our attention-based architecture was effective on a smaller scale, Nexus was designed to answer a critical question: how does this architecture perform when trained on a large, complex, and diverse dataset?
To this end, we took the robust attention-based architecture of Focalis, with its EfficientNetV2B3 backbone and Bi-LSTM spatial encoder, and trained it on the large-scale MS COCO 2017 dataset. MS COCO, with its  118k training images, is significantly larger and more complex than Flickr8k. Its images often contain multiple objects in intricate relationships, and it provides five reference captions per image, exposing the model to a much richer range of linguistic expression. The goal of Nexus was therefore to unlock the full capabilities of our architecture by leveraging this data. We hypothesized that the attention mechanism would learn more nuanced and accurate alignments, the visual encoder would learn more generalizable features, and the language decoder would develop a more sophisticated command of grammar and vocabulary. In essence, Nexus serves as the final validation of our iterative design process, demonstrating its effectiveness in a standard benchmark setting.
@@ 3.1 Experimental Setup
Datasets: Our initial models (Genesis-Focalis) were trained on the Flickr8k dataset, which contains approximately 8,000 images. Our final model, Nexus, was trained on the MS COCO 2017 dataset, using the official split of  118k training images and 5k validation images.
Training: The Nexus model was trained using the Adam optimizer [16] with an initial learning rate of $1\times 10^{-4}$ and ‘clipnorm=1.0‘. We employed a ‘ReduceLROnPlateau‘ scheduler (patience=1) and label smoothing ($\epsilon=0.1$) with the ‘CategoricalCrossentropy‘ loss function. Training was performed using a ”teacher forcing” strategy and continued until an ‘EarlyStopping‘ callback (patience=3) halted the process after 25 epochs.
Evaluation Metrics: We evaluate model performance using BLEU [13] and METEOR [14]. For inference, we use Beam Search with a beam width of $K=7$. As an independent researcher with limited computational resources, the widely-used CIDEr metric was not computed; however, BLEU and METEOR provide a robust evaluation consistent with foundational benchmarks and allow for clear comparison across our iterative models.
@@ 3.2 Result Analysis
The performance of each model, shown in Table 1, clearly illustrates the impact of our iterative architectural improvements and the critical role of data scaling.
Table 1: Performance Progression Across All Models. Models 1-4 are evaluated on Flickr8k; Nexus is evaluated on MS COCO.
Model Dataset BLEU-4 METEOR
Genesis Flickr8k 14.4 26.7
Contexta Flickr8k 17.2 36.1
Clarity Flickr8k 9.1 17.5
Focalis Flickr8k 20.2 38.1
Nexus MS COCO 31.4 47.4
@@
The Power of Scale: The most dramatic performance leap occurred with the Nexus model. By training our best architecture on the large-scale MS COCO dataset, we unlocked its true potential. The BLEU-4 score jumped from 20.2 to 31.4, an improvement of over 55%. This demonstrates that advanced architectures require large, diverse datasets to learn robust, generalizable representations. Our final model surpasses several foundational benchmarks, as shown in Table 4.
Qualitative Improvements: The impact of our architectural choices is most strikingly revealed when testing the models on challenging images from the MS COCO dataset, as shown in Table 2. This comparison exposes the combined limitations of both architecture and training data, demonstrating how the information bottleneck in the non-attention models forces them to ”hallucinate” when faced with unfamiliar scenes.
@@
Table 2: Qualitative Comparison of Sample Captions
Model Generated Caption
Image: A giraffe standing in a field (click to see)
Genesis A brown and white dog is running through the grass.
Clarity A dog is running through a grassy field.
Nexus A giraffe standing on top of a lush green field.
Image: Kites flying in the sky (click to see)
Genesis A group of people are playing in the water.
Clarity A group of people are walking on the beach.
Nexus There are many kites flying in the sky.
Image: People sitting at a table (click to see)
Genesis A group of people are standing in front of a crowd of people.
Clarity A group of people are sitting at a restaurant.
Nexus Several people sitting around a table with plates of food.
Training Dynamics and Model Selection
@@
The training process for Nexus revealed a divergence between the validation loss (our optimization target) and the final evaluation metrics. As shown in Figure 4, val_loss continued to improve until Epoch 25. However, an evaluation of checkpoints on a random 15-image sample (Table 3) showed that peak performance on metrics like F1-Score and BLEU-4 occurred earlier, around Epoch 13, where val_loss and train_loss first converged. After this point, the model began to over-optimize on the cross-entropy loss at the expense of caption quality. This finding is consistent with observations from prior work [6], where a breakdown in correlation between log-likelihood and BLEU was also noted. This analysis led us to select the checkpoint from Epoch 13 as our champion model for final evaluation and reporting.
@@
Table 3: Metric Comparison of Nexus Checkpoints on a Sample Set
Metric Epoch 10 Epoch 13 Epoch 25
BLEU-1 0.8352 0.7979 0.6066
BLEU-2 0.6961 0.6962 0.4643
BLEU-3 0.5428 0.5661 0.3294
BLEU-4 0.4192 0.4650 0.1856
METEOR 0.5557 0.5869 0.5878
Precision 0.8571 0.8367 0.8495
Recall 0.7372 0.7750 0.6049
F1-Score 0.7927 0.8047 0.6639
@@
Our analysis of the training dynamics highlights a significant divergence between the optimization objective (validation loss) and the task’s primary evaluation metrics (e.g., BLEU, METEOR). This difference arises because the CategoricalCrossentropy loss function serves as an imperfect proxy for sentence-level quality. While it effectively guides the model in next-word prediction, it does not explicitly reward n-gram overlap or semantic coherence across an entire generated sequence. We observe that after reaching a point of convergence around Epoch 13, further training leads to overfitting on this proxy objective. This results in the model favoring high-probability, often repetitive, phrases that continue to reduce loss but degrade overall caption quality, as reflected by decreasing BLEU scores. This misalignment between log-likelihood maximization and human-centric evaluation metrics is a recognized challenge in sequence generation, as previously discussed by Xu et al. [6].
@@
Consequently, although the model at Epoch 25 achieves the minimum validation loss, the checkpoint from Epoch 13 represents the optimal model for our task. It achieves the highest scores on metrics like BLEU-4 and METEOR, which are more closely aligned with human judgments of quality. This finding demonstrates the importance of using a comprehensive suite of task-specific metrics for model selection, rather than relying solely on the training loss.

### v2

# When Better Eyes Lead to Blindness: A Diagnostic Study of the Information Bottleneck in CNN-LSTM Image Captioning Models

@@ Abstract
Image captioning, situated at the intersection of computer vision and natural language processing, requires a sophisticated understanding of both visual scenes and linguistic structure. While modern approaches are dominated by large-scale Transformer architectures, this paper documents a systematic, iterative development of foundational image captioning models, progressing from a simple CNN-LSTM encoder-decoder to a competitive attention-based system. This paper presents a series of five models, beginning with Genesis and concluding with Nexus, an advanced model featuring an EfficientNetV2B3 backbone and a dynamic attention mechanism. The experiments chart the impact of architectural enhancements and demonstrate a key finding within the classic CNN-LSTM paradigm: merely upgrading the visual backbone without a corresponding attention mechanism can degrade performance, as the single-vector bottleneck cannot transmit the richer visual detail. This insight validates the architectural shift to attention. Trained on the MS COCO 2017 dataset, the final model, Nexus, achieves a BLEU-4 score of 31.4, surpassing several foundational benchmarks and validating the iterative design process. This work provides a clear, replicable blueprint for understanding the core architectural principles that underpin modern vision-language tasks.
keywords:
Image Captioning, Attention Mechanism, Information Bottleneck, Encoder-Decoder, CNN, RNN, LSTM, Spatial Encoder, MS COCO
terms: Deep Learning, Computer Vision, Natural Language Processing, Multimodal Learning
@@ 1 Introduction
In this paper, we present the results of a structured, iterative research process aimed at building and refining an image captioning model. The methodology is rooted in progressive enhancement, starting with a simple baseline and methodically introducing more complex components to address specific limitations. This paper documents the development of five distinct models, named to reflect their role in this evolution: Genesis, Contexta, Clarity, Focalis, and Nexus.
This journey provides valuable insights into the interplay between visual perception, language modeling, and attention. This paper demonstrates empirically how each architectural modification impacts performance, including a counter-intuitive but instructive finding where a more powerful visual encoder led to a decrease in caption quality, thereby motivating the shift to an attention-based design. This work also serves as a pedagogical guide for students and practitioners, offering a clear, replicable framework to understand the architectural trade-offs that shaped modern vision-language models. The final model, Nexus, leverages the lessons learned to achieve results that are competitive with and, on several key metrics surpass, foundational benchmarks in the field.
@@ 2 Related Work
The field of image captioning has evolved significantly since its inception. This section reviews the foundational encoder–decoder paradigms upon which this work builds, as well as the modern Transformer-based architectures that now define the state of the art.
@@ 2.1 Foundational Encoder-Decoder Architectures
Early work established the now-classic encoder-decoder architecture. Vinyals et al. [5] in their ”Show and Tell” model, successfully repurposed techniques from machine translation, using a pre-trained CNN like InceptionV3 as an encoder and an LSTM network as a decoder. The initial model, Genesis, is a direct implementation of this foundational paradigm.
The primary limitation of this approach was the information bottleneck created by compressing the entire image into a single static context vector. The breakthrough came with the introduction of visual attention mechanisms by Xu et al. [6] in ”Show, Attend and Tell”. Their model allowed the decoder to dynamically ”look” at different parts of the image at each step of the caption generation process. This concept of soft attention is central to the advanced models, Focalis and Nexus, and directly addresses the static bottleneck problem this paper investigates. Further refinements on this grid-based attention include the ”Bottom-Up and Top-Down Attention” model by Anderson et al. [7], which first identifies salient object regions and then applies top-down attention.
@@ 2.2 The Transformer Revolution and Modern Approaches
While attention-based Transformers are dominant, research into non-attention or attention-free architectures continues. For example, RWKV [24], a type of Recurrent Neural Network, has been adapted for multimodal tasks, offering a linear-time complexity alternative to the quadratic complexity of standard self-attention. This demonstrates that architectural diversity persists. Unlike the CNN-LSTM models with global average pooling that this paper investigates, architectures like RWKV or vision-specific designs like U-Nets [4] use different feature processing methods (recurrent states or skip connections) that may mitigate the specific single-vector bottleneck this paper identifies.
3 An Iterative, Diagnostic Approach to Model Design
The methodology is rooted in a structured, iterative research process designed to build and refine an image captioning model. This approach provides a crucial, hands-on analysis of the architectural trade-offs, particularly the attention-based resolution of the static information bottleneck that motivated the field’s transition towards more complex systems. This paper documents the development of five distinct models, progressively introducing more complex components to address specific limitations and test clear hypotheses.
@@ 3.1 Model 1: Genesis
The Genesis model serves as the foundational baseline, embodying the classic encoder-decoder paradigm that set the stage for modern image captioning. The architecture is intentionally simple to establish a clear performance benchmark. The encoder’s role is to ”see” the image, for which this paper employs the InceptionV3 architecture pre-trained on ImageNet. This choice provides a strong set of visual features learned from a massive dataset. The final convolutional feature map from InceptionV3 is passed through a Global Average Pooling layer, which condenses the entire spatial grid into a single, fixed-length 2048-dimension feature vector.
@@ 3.2 Model 2: Contexta
While Genesis established the visual baseline, its primary limitation was its decoder’s simplistic understanding of linguistic structure. A standard unidirectional LSTM only considers past context (the words it has already generated) when predicting the next word. To address this, this paper developed Contexta, a model designed to test the hypothesis that a richer language model could improve caption quality, even with the same static visual input.
@@ 3.3 Model 3: Clarity
The next step was to test a straightforward hypothesis: would a more powerful visual feature extractor directly lead to better captions? To investigate this, this paper designed the Clarity model. This paper replaces the InceptionV3 backbone with a state-of-the-art EfficientNetV2B3, known for its superior accuracy. Critically, this paper retained the rest of the Contexta architecture, including the global average pooling encoder and the Bi-LSTM decoder.
@@ 3.4 Model 4: Focalis
The performance degradation of Clarity provided a critical diagnostic insight: a more powerful visual feature extractor is useless if its output cannot be effectively utilized by the decoder. This confirmed the hypothesis that the single, static feature vector was acting as a severe information bottleneck. Although EfficientNetV2B3 was providing a richer, more detailed map of the image, global average pooling collapses this entire map into a single vector, irretrievably discarding the spatial information that defines ”where” objects are. To solve this, this paper designed Focalis, which represents a fundamental architectural pivot from a static to a dynamic vision-language interface.
The core innovation in Focalis is the introduction of a Luong-style attention mechanism, which completely removes the global average pooling layer. Instead of forcing the model to ”remember” the entire image from a single vector, This paper allows it to selectively ”look” at different regions of the image at each step of the caption generation process. This enables the decoder to ground specific words in specific spatial locations, for example, focusing on the region containing a ”dog” when generating that word, and then shifting focus to the ”ball” when describing what the dog is chasing.
@@
  • Feature Grid: The EfficientNetV2B3 encoder outputs a (10×10, 1536) feature grid, preserving the spatial layout of the image. After the spatial encoder (described below), this becomes a set of 100 context-aware feature vectors, which serve as the ”values” for the attention mechanism.
@@
Crucially, before the attention mechanism is applied, this paper introduce the key architectural refinement: a Bidirectional LSTM as a spatial encoder. Unlike standard attention mechanisms that treat each image patch independently, this Bi-LSTM processes the 100 patch vectors as a sequence. By reading the sequence of patches forwards and backwards, it enriches each patch’s vector with information about its neighbors. For instance, the vector for a patch containing a ”face” becomes aware that it is spatially adjacent to a patch containing ”hair.” This creates a more contextually rich set of feature vectors over which the decoder can attend.
@@
This complete architecture directly addresses the limitations observed in the previous models. By providing the decoder with dynamic, selective access to context-aware spatial features, this paper hypothesized that Focalis would not only resolve the performance degradation seen in Clarity but would significantly surpass all prior models. As the results show, this proved to be the case, marking the critical turning point in the development cycle.
@@ 3.5 Model 5: Nexus
The Nexus model does not introduce a new architecture; rather, it represents the culmination of the research by testing the scalability and full potential of the best design. While Focalis proved that the attention-based architecture was effective on a smaller scale, Nexus was designed to answer a critical question: how does this architecture perform when trained on a large, complex, and diverse dataset?
To this end, this paper took the robust attention-based architecture of Focalis, with its EfficientNetV2B3 backbone and Bi-LSTM spatial encoder, and trained it on the large-scale MS COCO 2017 dataset. MS COCO, with its  118k training images, is significantly larger and more complex than Flickr8k. Its images often contain multiple objects in intricate relationships, and it provides five reference captions per image, exposing the model to a much richer range of linguistic expression. The goal of Nexus was therefore to unlock the full capabilities of the architecture by leveraging this data. This paper hypothesized that the attention mechanism would learn more nuanced and accurate alignments, the visual encoder would learn more generalizable features, and the language decoder would develop a more sophisticated command of grammar and vocabulary. In essence, Nexus serves as the final validation of the iterative design process, demonstrating its effectiveness in a standard benchmark setting.
@@ 4.1 Experimental Setup
Datasets: The initial models (Genesis-Focalis) were trained on the Flickr8k dataset, which contains approximately 8,000 images. The final model, Nexus, was trained on the MS COCO 2017 dataset, using the official split of  118k training images and 5k validation images.
Training: The Nexus model was trained using the Adam optimizer [16] with an initial learning rate of $1\times 10^{-4}$ and ‘clipnorm=1.0‘. This paper employed a ‘ReduceLROnPlateau‘ scheduler (patience=1) and label smoothing ($\epsilon=0.1$) with the ‘CategoricalCrossentropy‘ loss function. Training was performed using a ”teacher forcing” strategy and continued until an ‘EarlyStopping‘ callback (patience=3) halted the process after 25 epochs.
Evaluation Metrics: this paper evaluate model performance using BLEU [13] and METEOR [14]. For inference, this paper use Beam Search with a beam width of $K=7$. As an independent researcher with limited computational resources, the widely-used CIDEr metric was not computed; however, BLEU and METEOR provide a robust evaluation consistent with foundational benchmarks and allow for clear comparison across the iterative models.
@@ 4.2 Result Analysis
The performance of each model, shown in Table 1, clearly illustrates the impact of the iterative architectural improvements and the critical role of data scaling.
[…]
Performance Progression Across All Models. Models 1-4 are evaluated on Flickr8k; Nexus is evaluated on MS COCO. Model Dataset BLEU-4 METEOR Genesis Flickr8k 14.4 26.7 Contexta Flickr8k 17.2 36.1 Clarity Flickr8k 9.1 17.5 Focalis Flickr8k 20.2 38.1 Nexus MS COCO 31.4 47.4
@@
The Power of Scale: The most dramatic performance leap occurred with the Nexus model. By training the best architecture on the large-scale MS COCO dataset, this paper unlocked its true potential. The BLEU-4 score jumped from 20.2 to 31.4, an improvement of over 55%. This demonstrates that advanced architectures require large, diverse datasets to learn robust, generalizable representations. The final model surpasses several foundational benchmarks, as shown in Table 4.
Qualitative Improvements: The impact of architectural choices is most strikingly revealed when testing the models on challenging images from the MS COCO dataset, as shown in Table 2. This comparison exposes the combined limitations of both architecture and training data, demonstrating how the information bottleneck in the non-attention models forces them to ”hallucinate” when faced with unfamiliar scenes.
@@
[…]
Qualitative Comparison of Sample Captions. Model Generated Caption Image: A giraffe standing in a field (see Appendix Fig. 6) Genesis A brown and white dog is running through the grass. Clarity A dog is running through a grassy field. Nexus A giraffe standing on top of a lush green field. Image: Kites flying in the sky (see Appendix Fig. 7) Genesis A group of people are playing in the water. Clarity A group of people are walking on the beach. Nexus There are many kites flying in the sky. Image: People sitting at a table (see Appendix Fig. 8) Genesis A group of people are standing in front of a crowd of people. Clarity A group of people are sitting at a restaurant. Nexus Several people sitting around a table with plates of food.
@@
Training Dynamics and Model Selection The training process for Nexus revealed a divergence between the validation loss (the optimization target) and the final evaluation metrics. As shown in Figure 4, val_loss continued to improve until Epoch 25. However, an evaluation of checkpoints on a random 15-image sample (Table 3) showed that peak performance on metrics like F1-Score and BLEU-4 occurred earlier, around Epoch 13, where val_loss and train_loss first converged. After this point, the model began to over-optimize on the cross-entropy loss at the expense of caption quality. This finding is consistent with observations from prior work [6], where a breakdown in correlation between log-likelihood and BLEU was also noted. This analysis led us to select the checkpoint from Epoch 13 as the champion model for final evaluation and reporting.
[…]
@@
Metric Comparison of Nexus Checkpoints on a Sample Set. Metric Epoch 10 Epoch 13 Epoch 25 BLEU-1 0.8352 0.7979 0.6066 BLEU-2 0.6961 0.6962 0.4643 BLEU-3 0.5428 0.5661 0.3294 BLEU-4 0.4192 0.4650 0.1856 METEOR 0.5557 0.5869 0.5878 Precision 0.8571 0.8367 0.8495 Recall 0.7372 0.7750 0.6049 F1-Score 0.7927 0.8047 0.6639
@@
The analysis of the training dynamics highlights a significant divergence between the optimization objective (validation loss) and the task’s primary evaluation metrics (e.g., BLEU, METEOR). This difference arises because the CategoricalCrossentropy loss function serves as an imperfect proxy for sentence-level quality. While it effectively guides the model in next-word prediction, it does not explicitly reward n-gram overlap or semantic coherence across an entire generated sequence. This paper observe that after reaching a point of convergence around Epoch 13, further training leads to overfitting on this proxy objective. This results in the model favoring high-probability, often repetitive, phrases that continue to reduce loss but degrade overall caption quality, as reflected by decreasing BLEU scores. This misalignment between log-likelihood maximization and human-centric evaluation metrics is a recognized challenge in sequence generation, as previously discussed by Xu et al. [6].
@@
Consequently, although the model at Epoch 25 achieves the minimum validation loss, the checkpoint from Epoch 13 represents the optimal model for the task. It achieves the highest scores on metrics like BLEU-4 and METEOR, which are more closely aligned with human judgments of quality. This finding demonstrates the importance of using a comprehensive suite of task-specific metrics for model selection, rather than relying solely on the training loss.